# DAGLFNet: Deep Attention-Guided Global–Local Feature Fusion for Pseudo-Image Point Cloud Segmentation

### v1

@@
DAGLFNet: Deep Attention-Guided Global–Local Feature Fusion for Pseudo-Image Point Cloud Segmentation
@@
Environmental perception systems play a critical role in high-precision mapping and autonomous navigation, with LiDAR serving as a core sensor that provides accurate 3D point cloud data. How to efficiently process unstructured point clouds while extracting structured semantic information remains a significant challenge, and in recent years, numerous pseudo-image-based representation methods have emerged to achieve a balance between efficiency and performance. However, they often overlook the structural and semantic details of point clouds, resulting in limited feature fusion and discriminability. In this work, we propose DAGLFNet, a pseudo-image-based semantic segmentation framework designed to extract discriminative features. First, the Global-Local Feature Fusion Encoding module is used to enhance the correlation among local features within a set and capture global contextual information. Second, the Multi-Branch Feature Extraction network is employed to capture more neighborhood information and enhance the discriminability of contour features. Finally, a Feature Fusion via Deep Feature-guided Attention mechanism is introduced to improve the precision of cross-channel feature fusion. Experimental evaluations show that DAGLFNet achieves 69.83% and 78.65% on the validation sets of SemanticKITTI and nuScenes, respectively. The method balances high performance with real-time capability, demonstrating great potential for LiDAR-based real-time applications.
@@ 1 Introduction
Semantic segmentation has emerged as a cornerstone technology for three-dimensional environmental perception, enabling dense semantic annotation and feature learning from LiDAR point clouds [1, 2, 3]. By providing structured interpretations of complex environments, this capability is fundamental to applications such as robotics and autonomous driving. Modern LiDAR sensors capture tens of millions of points per second, allowing an unprecedentedly detailed representation of the surrounding 3D structure [4, 5]. The central challenge, however, lies in devising effective strategies to process inherently unstructured and unordered point cloud data in order to extract discriminative features for reliable perception.
@@
Current LiDAR segmentation techniques adopt several distinct paradigms for handling point cloud data. Point-based methods process raw point sets directly, enabling dense feature interaction but relying on computationally intensive neighborhood searches to capture local geometric structures [6, 7, 8]. This leads to considerable resource demands and constrains their scalability in large outdoor environments. Voxel-based methods discretize point clouds into regular volumetric grids and apply sparse convolutions for feature extraction [9, 10, 11, 12]. Yet, the cubic growth of memory consumption with resolution makes high-resolution voxel grids prohibitively costly to construct and process. Hybrid strategies that integrate features from multiple representational domains have shown improved predictive accuracy [13]. Nevertheless, the heavy computational burden and issues of system robustness continue to pose major challenges for real-world deployment.
@@
Recently, range-image-based point cloud semantic segmentation has attracted increasing attention as it offers a favorable balance between computational efficiency and segmentation accuracy [3, 14, 15, 16]. By projecting irregular and complex LiDAR point clouds into structured two-dimensional representations such as range images [17, 18], this approach introduces several notable advantages. The regular image grid greatly simplifies neighborhood queries and local feature aggregation, while circumventing costly and time-consuming three-dimensional operations, and achieves promising performance, as shown in the Figure 1(b). However, this projection inevitably introduces structural distortions, causing blurred boundaries and semantic ambiguities that undermine the fidelity of structural representations, as shown in Figure 1(a). The principal challenge, therefore, lies in compensating for the loss of three-dimensional geometric information during projection, particularly in distant and inherently sparse regions.
@@
A fundamental challenge of pseudo-image-based methods lies in the partitioning of point clouds, where points from distinct semantic categories may be assigned to the same subregion. Such inconsistent grouping undermines the stability of local feature encoding, causing the pseudo-image to lose critical information regarding intra-set geometric and semantic relationships, and thereby diminishing the expressive value of the segmented regions. Furthermore, when features of these local point sets are projected into a two-dimensional image, high-dimensional geometric details are compressed into abstract 2D representations, leading to blurred boundaries and loss of fine-grained structural information, which exacerbates semantic ambiguity and feature degradation. In addition, the coarse fusion of 2D image features with the original point cloud representations can introduce redundant or conflicting information, limiting the discriminative capacity and fidelity of point-wise features. These issues are particularly pronounced in distant, sparsely populated, or occluded regions, where projected subregions often contain insufficient valid points, resulting in a marked decline in the structural fidelity of the pseudo-image representation.
To address the aforementioned challenges, we propose a novel framework, DAGLFNet, which integrates global–local feature aggregation, multi-branch feature extraction, and an attention-driven fusion mechanism. This design achieves a favorable balance between accuracy and computational efficiency in pseudo-image-based point cloud semantic segmentation. Specifically, local point sets generated through sub-region partitioning based on azimuth and laser-beam segmentation often exhibit significant internal geometric variability, which undermines the stability of local feature representations. To address this challenge and enhance the consistency of local geometric structures, we propose a Global-Local Feature Fusion Encoding (GL-FFE) module, capable of simultaneously capturing global contextual dependencies and fine-grained local geometric relationships. During the mapping of point cloud subsets into image representations, boundary features are often blurred and subject to interference from adjacent regions. To overcome this limitation, we propose a Multi-Branch Feature Extraction network (MB-FE), designed to expand the receptive field and enhance the discriminative representation of boundary features. In addition, we introduce a Feature Fusion via Deep Feature-guided Attention (FFDFA) strategy during the feature integration stage, which explicitly leverages distance information as a weighting constraint to enhance the precision of cross-channel feature fusion.
@@
    We introduce a novel network architecture, DAGLFNet, for semantic segmentation of LiDAR point clouds. Within this framework, geometric features of the point cloud are tightly integrated with two-dimensional pseudo-image representations, enabling efficient processing of unstructured and unordered point data while fully capturing discriminative point-wise features.
@@
    We propose a comprehensive feature enhancement strategy, comprising three key modules: (i) a GL-FFE module to capture long-range dependencies and stabilize local geometric representations; (ii) a MB-FE network to expand the receptive field and strengthen boundary feature expression; and (iii) a FFDFA mechanism that leverages distance-aware weighting to improve inter-channel feature integration.
@@
    Extensive experiments on two widely adopted LiDAR segmentation benchmarks demonstrate the superiority of DAGLFNet. Specifically, DAGLFNet achieves mean Intersection-over-Union (mIoU) scores of 69.8% and 78.7% on the validation sets of SemanticKITTI and nuScenes, respectively. Moreover, the framework can be successfully deployed on embedded platforms, enabling real-time semantic segmentation.
@@ 2 Related Work
2.1 LiDAR Point cloud segmentation
Point cloud semantic segmentation represents a cornerstone task in 3D perception [19, 20], aiming to assign precise semantic labels to individual points within a scene. By transforming raw and unstructured LiDAR data into dense, semantically enriched representations, this process enables a structured and fine-grained understanding of complex environments [21]. Among the diverse strategies developed for LiDAR point cloud semantic segmentation, point-based, voxel-based, and multi-modal data fusion approaches have emerged as three representative paradigms.
@@ 2.1 LiDAR Point cloud segmentation
Point-based methods operate directly on raw point clouds, preserving the full spatial geometric information of the scene. PointNet [8] pioneered the application of multi-layer perceptrons (MLPs) [22] to extract global feature correlations, inspiring subsequent architectures that integrate point convolution [23, 24] to capture local spatial patterns, graph convolution [25, 26] to model neighborhood relationships, and attention mechanisms [27, 28] to focus on salient features. These methods improved the ability of point cloud networks in local feature extraction and geometric structure preservation. However, while point convolution enhanced local pattern recognition, graph convolution improved neighborhood modeling, and attention mechanisms highlighted important features, all these approaches still suffered from high computational complexity when processing large-scale point clouds due to their dense operations and lack of efficient downsampling strategies. While RandLA-Net [6] mitigates computational demands through random sampling, point-based approaches remain challenged by the efficient processing of large-scale point clouds, limiting their applicability in real-time scenarios.
To impose structure on inherently unorganized point cloud data, voxel-based approaches discretize the point cloud into regular three-dimensional grids and employ 3D convolutional neural networks (3D CNNs) to extract hierarchical geometric features. Nevertheless, the intrinsic sparsity of point clouds renders voxel grids largely empty. OctNet [29] employs hierarchical octree decomposition to handle sparse point clouds, but still faces computational limitations in large-scale scenes. MinkNet [12] leverages Minkowski convolution operations to better capture geometric relationships in sparse voxel grids. Cylinder3D [11] introduces cylindrical voxelization specifically designed for LiDAR data, improving feature representation in cylindrical coordinate systems. However, these methods collectively face challenges in balancing computational efficiency with detailed feature extraction, as the voxel count in extensive LiDAR scenes still increases rapidly, incurring significant memory and computational overhead that limits their applicability in real-time scenarios.
@@
Recognizing the limitations of single data sources in fully capturing point cloud information, researchers have developed multi-modal fusion strategies to exploit the complementary strengths of diverse inputs. UniSeg [13] fuses voxel, view, and image features to fully utilize multi-modal semantic information. However, in practical applications, the heterogeneity of data sources in coordinate systems, resolution, and temporal synchronization makes precise feature alignment a significant challenge.
2.2 Semantic Segmentation Based on Range Images
@@
Recently, researchers have proposed a range of methodologies to optimize range-image-based semantic segmentation, aiming to enhance its precision [30, 14, 15, 16]. RangeNet++ [30] projects LiDAR point clouds onto spherical range images and applies convolutional neural networks for semantic segmentation, effectively reformulating the task into a structured two-dimensional representation. SalsaNext [31] and CENet [17] leverage enhanced contextual modeling to effectively mitigate boundary blurring and detail degradation inherent to pseudo-image-based methodologies. RangeFormer [16] partitions the entire LiDAR scan into multiple view-specific subsets and incorporates a Transformer-based architecture to capture long-range dependencies. Although these methods demonstrate reliable overall performance, they operate solely on points projected onto the image, struggle to resolve conflicts arising from multiple points mapping to the same location, overlook occluded points, and fail to preserve the full three-dimensional structure. To address this problem, FRNet [14] integrates point cloud and 2D features, preserving the fidelity of three-dimensional geometric information while leveraging the computational efficiency of two-dimensional convolutions. FARVNet [15] incorporates a reflectivity reconstruction module, amplifying the contribution of reflectivity to feature representation and thereby enhancing model robustness. Although achieving strong overall performance, these methods overlook intra-subset feature correlations, limiting their ability to accurately recognize sparsely occluded regions at long distances.
Distinct from prior approaches, this work emphasizes the coherent modeling of intra-subset feature relationships while systematically accounting for the effects of spatial distance, highlighting a more holistic consideration of local and contextual dependencies.
@@
Figure 2: The proposed DAGLFNet framework consists of key components such as GL-FFE, MB-FE, and FFDFA, which are responsible for contextual and geometric feature extraction, boundary enhancement, and multi-scale feature integration, respectively. Multiple stacked DAGLFNet units continuously learn complex hierarchical features from the point cloud. The Fusion Head combines point-level and group-level features to predict the final output.
@@ 3 Methodology
To overcome the limitations of pseudo-image-based segmentation, including point projection conflicts and incomplete 3D structure preservation, we propose a unified framework that effectively integrates global and local features. Our framework consists of three complementary components: a GL-FF module for capturing contextual and local geometric dependencies, a MB-FE network for enhancing boundary representations, and an AD-FF strategy for precise integration of multi-scale information. This design balances computational efficiency with high-fidelity 3D feature representation, addressing the key challenges inherent to range-image-based point cloud segmentation. The architecture overview of the proposed DAGLFNet framework is depicted in Figure 2.
@@ 3.1 Problem Definition
Given a point cloud $P={\{p_{1},p_{2},..,p_{N}\}}$ acquired by the LiDAR sensor, our network aims to assign a unique semantic label $L$ to each point, taking the point coordinates and reflectivity $p=(x,y,z,I)$ as input. This process can be summarized as:
@@
where $\theta$ represents the learnable parameters within the network. As illustrated in Figure 2, DAGLFNet comprises four key steps: 1) a global–local point cloud feature encoder for point cloud representation; 2) an image feature encoder for extracting semantic features; 3) a point cloud feature fusion module guided by depth values via an attention mechanism; 4) a fusion head that integrates multi-level point cloud features to achieve precise semantic prediction.
@@ 3.2 Feature Encoder
Given the limited information available from point clouds, effective feature encoding is critical [32]. Unlike voxelization or regional partitioning [11, 30], we divide the point cloud into $M$ groups $P=\{\mathbb{P}_{1},\mathbb{P}_{2},...,\mathbb{P}_{M}\}$ as follows:
@@
where $h_{l}$ and $\phi_{l}$ refer to the $l$-th Velodyne sensor. We then compute the projection coordinates of points $(r,\alpha,\phi)$ as $u=((\alpha+\pi)2\pi)W$, $v=l$, where $(u,v)$ denotes the grid coordinate of a point in the range image with a resolution of $H\times W$.
We first obtain the virtual center of each group by applying average pooling over all points within the group, denoted as $p_{avg}$. Based on this, the initial features of each point comprising its 3D coordinates, reflectance, depth, and offsets relative to the $p_{avg}$, are represented as $\bar{p}_{j}\in\mathbb{R}^{1\times 10}$. The features for each point are given by:
$\bar{p}_{j}=([x,y,z,depth,I;p_{j}-p_{avg}])$ (3)
Collectively, the features of all points form the point cloud feature matrix $\bar{P}=\{\bar{p}_{1},\bar{p}_{2},...,\bar{p}_{j}\}\in{\mathbb{R}^{N\times 10}}$. Subsequently, MLP [22] is employed to extract point cloud features. Specifically, the per-point features are first encoded by an MLP, yielding per-point feature representation $\mathcal{F}_{p}^{0}$. Group-level features are then obtained by aggregating the features of all points within each group using both maximum and average pooling. The outputs of these two pooling operations are combined to form the final group-level representation. The outputs of these two pooling operations are fused and projected to yield the final group-level representation $\mathcal{F}_{g}^{0}$:
@@
where $\text{Flat}(\cdot):\mathbb{R}^{N\times C}\to\mathbb{R}^{H\times W\times C}$ denotes the function that maps point features onto the range image plane with resolution $(H,W)$. In our network, $\mathcal{F}_{p}^{0}$ and $\mathcal{F}_{g}^{0}$ serve as inputs, $\mathcal{F}_{g}^{0}$ encodes global point cloud information while integrating group-level features, forming a comprehensive representation that captures both global contextual dependencies and local geometric structures, and enhances feature expressiveness.

### v2

# DAGLFNet: Deep Feature Attention Guided Global and Local Feature Fusion for Pseudo-Image Point Cloud Segmentation

@@
DAGLFNet: Deep Feature Attention Guided Global and Local Feature Fusion for Pseudo-Image Point Cloud Segmentation
@@
Environmental perception systems are crucial for high-precision mapping and autonomous navigation, with LiDAR serving as a core sensor providing accurate 3D point cloud data. Efficiently processing unstructured point clouds while extracting structured semantic information remains a significant challenge. In recent years, numerous pseudo-image-based representation methods have emerged to balance efficiency and performance. However, these methods often overlook the structural and semantic details of point clouds, leading to limited feature fusion and discriminability. This work proposes DAGLFNet, a pseudo-image-based semantic segmentation framework designed to extract discriminative features. It incorporates three key components: first, a Global-Local Feature Fusion Encoding (GL-FFE) module to enhance intra-set local feature correlation and capture global contextual information; second, a Multi-Branch Feature Extraction (MB-FE) network to capture richer neighborhood information and improve the discriminability of contour features; and third, a Feature Fusion via Deep Feature-guided Attention (FFDFA) mechanism to refine cross-channel feature fusion precision. Experimental evaluations demonstrate that DAGLFNet achieves mean Intersection-over-Union (mIoU) scores of 69.9% and 78.7% on the validation sets of SemanticKITTI and nuScenes, respectively. The method achieves an excellent balance between accuracy and efficiency, demonstrating strong potential for LiDAR-based applications.
@@ 1 Introduction
Semantic segmentation has become a cornerstone technology for 3D environmental perception, enabling dense semantic annotation and feature learning from LiDAR point clouds [1, 2, 3]. By providing structured interpretations of complex environments, this capability is fundamental to applications like robotics and autonomous driving. Modern LiDAR sensors capture tens of millions of points per second, allowing an unprecedentedly detailed representation of the surrounding 3D structure [4, 5]. However, the central challenge lies in developing effective strategies to process inherently unstructured and unordered point cloud data to extract discriminative features for reliable perception.
@@
Current LiDAR segmentation techniques employ several distinct paradigms for handling point cloud data. Point-based methods process raw point sets directly, enabling dense feature interaction but relying on computationally intensive neighborhood searches to capture local geometric structures [6, 7, 8]. This results in significant resource demands, limiting scalability in large outdoor environments. Voxel-based methods discretize point clouds into regular volumetric grids and apply sparse convolutions for feature extraction [9, 10, 11, 12]. However, the cubic growth of memory consumption with resolution makes high-resolution voxel grids prohibitively expensive to construct and process. Hybrid strategies integrating features from multiple representational domains have shown improved predictive accuracy [13]. Nevertheless, their heavy computational burden and challenges in system robustness remain significant obstacles for further improvement.
@@
In recent years, range-view-based semantic segmentation of point clouds has garnered increasing attention due to its favorable balance between computational efficiency and segmentation accuracy [3, 30]. By projecting irregular and complex LiDAR point clouds onto structured two-dimensional representations (such as range images), this approach enables more effective organization and encoding of spatial information, extraction of local and global features [17, 18], and offers several significant advantages. The regular image grids greatly simplify neighborhood queries and local feature aggregation, while avoiding costly and time-consuming three-dimensional operations, and have achieved encouraging performance. However, this projection strategy inevitably leads to depth conflict phenomena, where multiple 3D points are mapped to the same 2D pixel coordinates. This causes subsequent projected points to overwrite previously mapped points, resulting in an irreversible loss of 3D spatial structural information.
Recent methods have addressed the inherent depth conflicts and 3D information loss in direct range image projection by converting point clouds into pseudo-images. These methods aggregate and encode local 3D point clusters into 2D grid representations and fuse point features with 2D features to preserve complete point cloud information [14, 15], the segmentation results are shown in Fig. 1(b). However, points from distinct semantic categories may be assigned to the same subregion during point cloud partitioning, causing inconsistent grouping that destabilizes local feature encoding and diminishes the expressive value of segmented regions. Furthermore, projecting features of these local point sets into 2D images compresses high-dimensional geometric details into abstract representations, leading to blurred boundaries and loss of fine-grained structural information. This exacerbates semantic ambiguity and feature degradation, as shown in Fig. 1(a). Additionally, coarse fusion of 2D image features with original point cloud representations can introduce redundant or conflicting information, limiting the discriminative capacity and fidelity of point-wise features. These issues are particularly pronounced in distant, sparse, or occluded regions, where projected subregions often contain insufficient valid points, resulting in a significant decline in the structural fidelity of the pseudo-image representation.
@@
To address these challenges, we propose DAGLFNet, a novel framework integrating global-local feature aggregation, multi-branch feature extraction, and a deep feature-guided fusion mechanism. This design achieves a favorable balance between accuracy and computational efficiency for pseudo-image-based point cloud semantic segmentation. Specifically, local point sets generated by azimuth and laser-beam-based sub-region partitioning often exhibit significant internal geometric variability, undermining local feature representation stability. To address this and enhance local geometric consistency, we propose a Global-Local Feature Fusion Encoding (GL-FFE) module capable of simultaneously capturing global contextual dependencies and fine-grained local geometric relationships. During the mapping of point cloud subsets to image representations, boundary features are often blurred and susceptible to interference from adjacent regions. To overcome this limitation, we propose a Multi-Branch Feature Extraction (MB-FE) network designed to expand the receptive field and enhance the discriminative representation of boundary features. Additionally, we introduce a Feature Fusion via Deep Feature-guided Attention (FFDFA) strategy during feature integration, which explicitly leverages distance information as a weighting constraint to refine cross-channel feature fusion precision.
@@
    We introduce DAGLFNet, a novel network architecture for LiDAR point cloud semantic segmentation. Within this framework, geometric features are tightly integrated with 2D pseudo-image representations, enabling efficient processing of unstructured point data while fully capturing discriminative point-wise features.
@@
    We propose a comprehensive feature enhancement strategy comprising three key modules: (i) a GL-FFE module to capture long-range dependencies and stabilize local geometric representations; (ii) a MB-FE network to expand the receptive field and strengthen boundary feature representation; and (iii) a FFDFA mechanism leveraging distance-aware weighting to improve inter-channel feature integration.
@@
    Extensive experiments on two widely adopted LiDAR segmentation benchmarks demonstrate the superiority of DAGLFNet. Specifically, it achieves mean Intersection-over-Union (mIoU) scores of 69.8% and 78.7% on the SemanticKITTI and nuScenes validation sets, respectively.
@@ 2 Related Work
2.1 Methodologies for LiDAR Point Cloud Semantic Segmentation
Point cloud semantic segmentation assigns semantic labels to individual points in LiDAR scenes, transforming unstructured data into structured environmental understanding [19, 20, 21]. This section focuses on three prominent paradigms: point-based, voxel-based, and multi-modal fusion methods.
@@ 2.1 Methodologies for LiDAR Point Cloud Semantic Segmentation
Point-based methods operate directly on raw point clouds, preserving the full spatial geometric information of the scene. PointNet [8] pioneered the application of multi-layer perceptrons (MLPs) [22] to extract global feature correlations, inspiring subsequent architectures that integrate point convolution [23, 24] to capture local spatial patterns, graph convolution [25, 26] to model neighborhood relationships, and attention mechanisms [27, 28] to focus on salient features. These methods enhance local feature extraction and geometric structure preservation capabilities. However, when processing large-scale point clouds, computational complexity remains high due to intensive computation and the lack of efficient downsampling strategies.
Voxel-based methods convert point clouds to regular 3D grids using 3D CNNs for geometric feature extraction, yet face computational burdens due to sparsity. OctNet [9] addresses this through octree decomposition to reduce redundant calculations, though still encounters bottlenecks in large-scale scenarios. To further optimize sparse grid processing, MinkNet [12] introduces Minkowski convolution operations that better capture geometric relationships in sparse voxel grids, achieving higher efficiency than traditional voxelization methods. Meanwhile, Cylinder3D [11] employs cylindrical voxelization specifically designed for LiDAR data, enabling better feature representation in cylindrical coordinates and particularly suitable for point clouds with strong structural characteristics in specific viewing angles. However, these methods face trade-offs between computational efficiency and detailed feature extraction, as rapidly increasing voxel counts in large LiDAR scenes impose significant computational overhead.
@@
In response to the inherent limitations of single data sources in comprehensively capturing point cloud information, multi-modal fusion strategies have been developed to leverage the complementary advantages of heterogeneous input data. UniSeg [13] fuses voxel, view, and image features to fully utilize multi-modal semantic information. However, in practical applications, the heterogeneity of data sources in coordinate systems, resolution, and temporal synchronization makes precise feature alignment a significant challenge.
2.2 2D Representation for Semantic Segmentation
@@
To achieve a balance between performance and efficiency in LiDAR point cloud semantic segmentation, researchers have proposed range image-based semantic segmentation methods [30, 14, 15, 16]. RangeNet++ [30] projects LiDAR point clouds onto spherical range images and applies convolutional neural networks for semantic segmentation, effectively reformulating the task into a structured two-dimensional representation. SalsaNext [31] and CENet [17] leverage enhanced contextual modeling to effectively mitigate boundary blurring and detail degradation inherent to pseudo-image-based methodologies. RangeFormer [16] partitions the entire LiDAR scan into multiple view-specific subsets and incorporates a Transformer-based architecture to capture long-range dependencies. Although these methods demonstrate reliable overall performance, they operate solely on points projected onto the image, struggle to resolve conflicts arising from multiple points mapping to the same location, overlook occluded points, and fail to preserve the full three-dimensional structure. To address this problem, FRNet [14] integrates point cloud and 2D features, preserving the fidelity of three-dimensional geometric information while leveraging the computational efficiency of two-dimensional convolutions. FARVNet [15] incorporates a reflectivity reconstruction module, amplifying the contribution of reflectivity to feature representation and thereby enhancing model robustness. Although achieving strong overall performance, these methods overlook intra-subset feature correlations, limiting their ability to accurately recognize sparsely occluded regions at long distances. Distinct from prior approaches, this work emphasizes the coherent modeling of intra-subset feature relationships while systematically accounting for the effects of spatial distance, highlighting a more holistic consideration of local and contextual dependencies.
@@
Figure 2: The proposed DAGLFNet framework consists of key components such as GL-FFE, MB-FE, FFDFA, and the Fusion Head, which are responsible for contextual and geometric feature extraction, boundary enhancement, multi-scale feature integration, and final prediction respectively. Multiple stacked DAGLFNet units continuously learn complex hierarchical features from the point cloud, with the Fusion Head combining point-level and group-level features to predict the final output.
@@ 3 Methodology
Building on the limitations identified in Section 2.2, we propose a unified framework that jointly optimizes 2D projection and 3D structural fidelity. Our framework consists of four complementary components: a GL-FFE module for capturing contextual and local geometric dependencies, a MB-FE network for enhancing boundary representations, a FFDFA strategy for precise integration of multi-scale information, and a fusion head to integrate low-level and high-level semantic information. The architecture overview of the proposed DAGLFNet framework is depicted in Fig. 2.
@@ 3.1 Problem Definition
Given a point cloud $P=\{p_{1},p_{2},\ldots,p_{N}\}$, where each point $p_{j}=(x_{j},y_{j},z_{j},I_{j})$ consists of its 3D coordinates $(x_{j},y_{j},z_{j})$ and reflectivity $I_{j}$ captured by the LiDAR sensor, our network aims to predict a semantic label $l_{i}\in L$ for each point $p_{j}$ through the network architecture $\mathcal{G}$. Formally, the network can be defined as:
@@
where $\theta$ represents the learnable parameters within the network. As illustrated in Fig. 2, DAGLFNet comprises four key steps: 1) a global–local point cloud feature encoder for point cloud representation; 2) an image feature encoder for extracting semantic features; 3) a point cloud feature fusion module guided by depth values via an attention mechanism; 4) a fusion head that integrates multi-level point cloud features to achieve precise semantic prediction.
@@ 3.2 Feature Encoder
Given the sparsity of point cloud data [32], the design of feature encoder is crucial for effectively extracting features and converting them into 2D representations. Unlike conventional voxelization or regional partitioning approaches [11, 30], we propose a grouping strategy based on laser beam characteristics to divide the point cloud into $M$ groups $P={\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{M}\}}$, as follows:
@@
where $h_{l}$ and $\phi_{l}$ denote the vertical offset and the elevation angle of the $l$-th laser beam, respectively. We then compute the projection coordinates of points $(r,\alpha,\phi)$ as $u=((\alpha+\pi)2\pi)W$, $v=l$, where $(u,v)$ denotes the grid coordinate of a point in the range image with a resolution of $H\times W$.
To comprehensively capture geometric and contextual information within each group, the virtual center $p_{avg}$ is obtained by applying average pooling over all points in the group. This center serves as a reference for encoding relative spatial relationships. The initial feature representation of each point, denoted as $\bar{p}_{j}\in\mathbb{R}^{1\times 10}$, is constructed by integrating multiple complementary information sources, including absolute 3D coordinates, reflectance intensity, depth (Euclidean distance to the sensor), and relative offsets to the group center $p_{avg}$. This multi-modal feature encoding allows the network to learn both global context and local geometric details. The feature vector for each point is defined as:
$\bar{p}_{j}=([x_{j},y_{j},z_{j},I_{j},depth_{j};p_{j}-p_{avg}])$ (3)
Collectively, the features of all points form the point cloud feature matrix $\bar{P}=\{\bar{p}_{1},\bar{p}_{2},...,\bar{p}_{j}\}\in{\mathbb{R}^{N\times 10}}$. Subsequently, an MLP [22] is employed to extract discriminative point cloud features. Each point feature $\bar{p}_{j}$ is first encoded by an MLP, yielding the initial per-point feature representation $\mathcal{F}_{p}^{0}\in\mathbb{R}^{N\times C}$. Although the point-level features $\mathcal{F}_{p}^{0}$ already encode the context of each point within the entire point cloud, they do not explicitly model the geometric relationships among points within each group. To address this, features within each group $\mathcal{P}_{M}$ are aggregated to obtain compact group-level representations. Specifically, max pooling and average pooling are applied to the point-level features to capture complementary structural patterns. The resulting features are concatenated and projected through a linear mapping followed by ReLU activation, yielding the final group-level features, $\mathcal{F}_{g}^{0}\in\mathbb{R}^{H\times W\times C}$:
@@
where $\text{Flat}(\cdot):\mathbb{R}^{N\times C}\to\mathbb{R}^{H\times W\times C}$ denotes the function that maps point features onto the range image plane with resolution $(H,W)$. $\mathcal{F}_{p}^{0}$ preserves individual point characteristics, while $\mathcal{F}_{g}^{0}$ incorporates global contextual understanding through aggregated group-level features, forming a complementary representation.